Solve the equation.
1 + 5x = -34

Answer by addingup(3677)

Always do the same to both sides:
1 + 5x = -34 subtract 1 to both sides
1 - 1 + 5x = -34 - 1
5x = -35 divide both sides by 5
5x/5 = -35/5 on the left, 5/5 cancel each other and you are left with x
On the right, divide -35/5
x = -7
.
Happy learning